OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80 A. an = -5 • (-4)n B. an = -5(-4)n-1 C. an = -4(-5)n-1 D. an = -5 • (4)n

OpenStudy (anonymous):

@GraceJaylene @deezgurls I got D.

OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1 And C here

OpenStudy (anonymous):

@Peter14

OpenStudy (anonymous):

@Frostbite HELP! plz :)

OpenStudy (anonymous):

@GraceJaylene !!!!! DDDD:

OpenStudy (amistre64):

a1 does not equall; -5*4(1)

OpenStudy (anonymous):

yeah could u explain this to me? :)

OpenStudy (amistre64):

since its geometric, we want to determine a common "r" such that a1 = -5 * r = 20 a2 = 20 * r = -80 a3 = -80 r = 4 is good and your options appear to be missing the ^ to indicate an exponent

OpenStudy (amistre64):

in any case an = a1 * r^(n-1) is the explicit formula for a geometric seq

OpenStudy (amistre64):

err, -4 is fine for r

OpenStudy (amistre64):

the recurrsive mirrors the setup that gets you to find "r" a1 = k * r = a2 * r = a3 *r = .....= an-1 * r = an * r = an+1*r = .....

OpenStudy (anonymous):

ohh

OpenStudy (amistre64):

a1 = -2 * r = 8 a2 = 8 * r = -32 a3 = -32 + r = .... an = r * an-1

OpenStudy (anonymous):

i see it :P

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thanks :P